OpenStudy (anonymous):

In simplified exponential notation, the expression x ³ • x -4 • x = _____

OpenStudy (michele_laino):

please we have \[{x^3} \cdot x = {x^4}\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so I can rewrite your expression as follows: \[{x^4} - 4x\]

OpenStudy (anonymous):

x3.x-4.x=

OpenStudy (anonymous):

so is 0?

OpenStudy (anonymous):

or 1/x ?

OpenStudy (michele_laino):

no, please it is: \[{x^3} \cdot x - 4 \cdot x = {x^4} - 4x\]

OpenStudy (anonymous):

do not understand i only have 3 choices 1/x 0 1

OpenStudy (anonymous):

?

OpenStudy (michele_laino):

is your original expression this? \[\left( {{x^3} \cdot x} \right) - \left( {4 \cdot x} \right)\]

OpenStudy (anonymous):

1/x?

OpenStudy (michele_laino):

please I aks help to another tutor @Abhisar

OpenStudy (anonymous):

ty

OpenStudy (michele_laino):

@Nnesha

OpenStudy (anonymous):

x3.x-4.x=

OpenStudy (abhisar):

\(\tt x^{-4}~can~be~written~as~\frac{1}{x^4}\)

OpenStudy (solomonzelman):

IS the initial problem, \(\large\color{slate}{ x^3\cdot x^{-4}\cdot x }\) ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^3\cdot x^{-4}\cdot x =x^{3-4+1}=x^{?} }\)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

what is the exponent here?

OpenStudy (solomonzelman):

what is your expoentn can you tell me?

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

and the prove for \(\large\color{slate}{ x^0 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{slate}{ \displaystyle \color{red}{x^{0}}~~\Rightarrow ~~x^{a-a}~~\Rightarrow ~~\frac{x^a}{x^a}~~\Rightarrow ~~\color{red}{1} }\)

OpenStudy (solomonzelman):

so what is your final answer ?

OpenStudy (anonymous):

1/x

OpenStudy (solomonzelman):

wait, you said the exponent is 0, \(\large\color{slate}{ x^3\cdot x^{-4} \cdot x~~~\Longrightarrow~~~ x^3\cdot x^{-4} \cdot x^1~~~\Longrightarrow~~~ x^{3 +(-4)+1}~~~\Longrightarrow~~~ x^0. }\)

OpenStudy (solomonzelman):

And as I showed \(\large\color{slate}{ x^0 }\) is equivalent to what number ?

OpenStudy (anonymous):

This is so confusing :(

Nnesha (nnesha):

\[2^0 = 1\] \[4^0 =1\] \[b^0 = 1\] when any number or variable have 0 exponent answer always going to be just 1

Nnesha (nnesha):

you said that x^0 right ??? that's what you got right ?? @Dyalis_love

Nnesha (nnesha):

are you there ???